OpenStudy (anonymous):

Ann had 45 fliers to post around town. Last week, she posted 2/3 of them. This week, she posted 4/5 of the remaining fliers. How many fliers has she still not posted? i got 6 3/4 but it said i was wrong :(

OpenStudy (anonymous):

24 is that right?

OpenStudy (anonymous):

So the first week she posted 45 (2/3) or 30 of them and had 15 left. This week she posted 15 (4/5) of them so had 12 left.

OpenStudy (anonymous):

So, 1/3 of the posters were not originally posted. Then, 1/5 of the remaining 1/3 were not posted. So, 1/5 of 1/3 is 1/15. 45/15 = 3.

OpenStudy (anonymous):

oops I meant 3 left

OpenStudy (anonymous):

THREE. :3

OpenStudy (anonymous):

12 is the right answer I am pretty sure

OpenStudy (anonymous):

three was right! :) i cant get this one either. There are 75 students in a speech contest. Yesterday, 1/5 of them gave their speeches. Today, 1/3 of the remaining students gave their speeches. How many students still haven't given their speeches?

OpenStudy (anonymous):

Again 75 total and 1/5 gave their speech -> 75 * 1/5 = 15 60 were left (75 - 15) 60 and 1/3 gave their speech -> 60 * 1/3 = 20 40 left (60 - 20)

OpenStudy (anonymous):

1/5 of them were omitted the first time. 1/3 of the 1/5 then gave speeches. 1/5 of 1/3 is 1/15, again. 75/15 = 5.

OpenStudy (anonymous):

** who have given their speech...then 75 - 5 = 7- students.

OpenStudy (anonymous):

*** ugh, stupid keyboard...70.

OpenStudy (anonymous):

40?

OpenStudy (anonymous):

it said 70 was wrong.

OpenStudy (anonymous):

Wow, I completely screwed that problem up..:P sorry, I was distracted xD

OpenStudy (anonymous):

awkward... looks like TBates beat you too it

OpenStudy (anonymous):

40 is what it should be.

OpenStudy (anonymous):

Stfu Chad.

OpenStudy (anonymous):

haha sorry...

OpenStudy (anonymous):

great then what would this one be?? There are 90 new houses being built in a neighborhood. Last month, 1/3 of them were sold. This month, 1/6 of the remaining houses were sold. How many houses are left to be sold?

OpenStudy (anonymous):

this one i'll try

OpenStudy (anonymous):

50!

OpenStudy (anonymous):

45..

OpenStudy (anonymous):

She needs more than to just be told the answer. 1/3 of the 90 houses were sold, leaving 60. 1/6 of the 60 were sold, leaving 50. That's how Chad got his answer.

OpenStudy (anonymous):

i got 50

OpenStudy (anonymous):

Sh*t.

OpenStudy (anonymous):

tbates am i right?

OpenStudy (anonymous):

You are!

OpenStudy (anonymous):

ok just because he has a giant green "pi" as his picture doesnt mean he is better than us at math...

OpenStudy (anonymous):

"You are!"

OpenStudy (anonymous):

lol he has just been right every time i've asked him. no offense!!! :)

OpenStudy (anonymous):

A bookstore had 64 copies of a magazine. Yesterday, it sold 1/4 of them. Today, it sold 3/4 of what remained. How many copies does the bookstore have left? i got 28 Am i correct???

OpenStudy (anonymous):

i think it should be 12

OpenStudy (anonymous):

It should be 12 like chad said

OpenStudy (anonymous):

thank you